OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours. : Age Ticket Price Under 3 years free 3 to 12 years $32 Over 12 years $44 On a certain day, the company took 82 people on whale watching trips. There were 54 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (54 + x)32 + (82 - (54 + x)) ⋅ 44 = C (54 - x)32 + (82 - 54) ⋅ 44 = C

OpenStudy (anonymous):

(82 + x)32 + (82 - (54 + x)) ⋅ 44 = C (82 - x)32 + (82 - 54) ⋅ 44 = C